OpenStudy (anonymous):

Prove tan(theta/2)+cot(theta/2) = 2csc(theta)

OpenStudy (anonymous):

\[ \tan(\theta/2)=\frac{\sin(\theta/2)}{\cos(\theta/2)} \]